hello,
this problem is driving me crazy. It is,in paranthesis, (3 sqrt 5)^5
Could you please show me how it is worked?
Thank you

Answer by stanbon(57387)


(3 sqrt 5)^5
= 3^5(sqrt5)^5
=243 (5)^(5/2)
=243(5^2)(5^(1/2))
=243(25)sqrt5
=6075sqrt5
Cheers,
Stan H.